Kamala had 5,026 grams of flour in a canister. She brought a 4,157 gram bag of flour. She poured some flour from the bag to the canister. As a result, the mass of the flour in the canister is now twice the mass of the flour left in the bag. How much flour is in the bag now?
Answer #1
5026 +(4157-x) = 2 (x)

5026 + 4157 = 3x

x = 3061 left in the bag

check
amount poured = 4157-3061 = 1096
amount in canister = 1096+5026 = 6122
6122/ 2 = 3061 Caramba !